jhonyy9 (jhonyy9):

- how is this possible ? 1×8 +1 =9 12×8 +2 =98 123×8 +3 =987 1234×8 +4 =9876 12345×8 +5 =98765 123456×8 +6 =987654 1234567×8 +7 =9876543 12345678×8 +8 =98765432 123456789×8 +9 =987654321

alones (alones):

ooo lalala a pyramid

OpenStudy (agent0smith):

Interesting.

jhonyy9 (jhonyy9):

can be this realized using different number in place of 8 or just in this way is possible ?

jhonyy9 (jhonyy9):

@TheSmartOne

OpenStudy (kainui):

I noticed we can write this, to take a specific example: \[1234560-123456+6=1111110\]\[123456*(10-1)+6=1111110\]\[123456*9+6=1111110\]\[123456*8+6=1111110-123456\]\[123456*8+6=987654\] The main two things I see going on are we have this relationship here which isn't too hard to see: \[123456+987654=1111110\] And the other relationship we have it \(9=10-1\) which is what allows us to put the numbers next to each other so that we get the digits to line up like 9-8=1, 8-7=1, 7-6=1, etc... Hopefully this helps in trying to find some other kinds of patterns like this!